OpenStudy (anonymous):

Find the equation of the circle with center that passes through point A. center (1, 3) and point A (4, -2) equaling to 0

OpenStudy (amistre64):

whats the equation of a generic circle?

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

Use the distance formula to find the radius

OpenStudy (amistre64):

good, now replace h with 1 and k with 3 for starters

OpenStudy (amistre64):

the radius "r" is just the distance from the center to A

OpenStudy (anonymous):

(x-1)^2+(y-3)^2=r^2 .....

OpenStudy (amistre64):

good so far; now for "r"

OpenStudy (amistre64):

do you recall the formula or any method to find the distance between 2 points?

OpenStudy (anonymous):

yeahhhhh

OpenStudy (amistre64):

then plug in the point for center and the point for A to determine the distance; thats your r then

OpenStudy (amistre64):

i get 3 over and 5 down sooo; sqrt(3^2+5^2) = sqrt(34) perhaps?

OpenStudy (amistre64):

and r^2 is then 34 (x-1)^2 + (y-3)^2 = 34 expand it all out and put all things to one side to zero it out

OpenStudy (anonymous):

wait im not sure how to zero that out..

OpenStudy (amistre64):

expand your ^2 terms

OpenStudy (amistre64):

whats (x-1)^2 ?

OpenStudy (anonymous):

x^2-1^2

OpenStudy (amistre64):

not quite, (x-1)(x-1) = x^2 -2x +1

OpenStudy (amistre64):

expand the (y-3)^2

OpenStudy (anonymous):

y^2-6^2+3

OpenStudy (amistre64):

.... youre really bad at that :) y-3 y-3 ---- y^2 -3y -3y+9 ---------- y^2 -6y + 9

OpenStudy (amistre64):

sooo: x^2 -2x +1 + y^2 -6y +9 = 34 combine the constants x^2 -2x + y^2 -6y +10 = 34 and -34 from each side x^2 -2x + y^2 -6y +10 = 34 -34 -34 --------------------------- x^2 -2x + y^2 -6y - 24 = 0

OpenStudy (anonymous):

gochaaa , thanks!

OpenStudy (amistre64):

good luck :)